OpenStudy (anonymous):

Solve 15^2x = 36. Round to the nearest ten-thousandth.

OpenStudy (anonymous):

@jim_thompson5910 you helped me with this ealier and i got it wrong

OpenStudy (anonymous):

A. 0.6616 C. 1.7509 D. 1.9091

OpenStudy (anonymous):

Is this \[15^{2x} = 36\]for the equation you are working with?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can start by taking the log of both sides and use the relationship that\[\ln a ^{b} = b \times \ln a\]

OpenStudy (anonymous):

whats with the b

OpenStudy (anonymous):

"b" is your "2x". "2x" will come off of the exponent and be a factor in front. I used "b" to show you a formula.

OpenStudy (anonymous):

im sorry but im way confused

OpenStudy (anonymous):

Can you get the exponent off of the following expression: ln [15^(2x)]. "2x" is your "b" in my example. It will go in front like the "b" went in front. Can you re-write this?

OpenStudy (anonymous):

0.6616

OpenStudy (anonymous):

@jim_thompson5910 i missed like all the graphs you helped me on

OpenStudy (anonymous):

In your work with logarithms, have you ever seen the formula \[\ln a ^{b} = b \times \ln a\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then part of your difficulty is unfamiliarity with logarithms. You would need a tutorial of some good length on the subject which is beyond the scope of answering this question in this forum.